OpenStudy (anonymous):

Help please... system with 3 variables 3x+2y-z=12 -4x+y-2z=4 x-3y+z= -4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what methods do you know

OpenStudy (anonymous):

like trying to eliminate 1 of the variables

jimthompson5910 (jim_thompson5910):

ok, if you add equations 1 and 3, what do you get

OpenStudy (anonymous):

4x-1y+8

jimthompson5910 (jim_thompson5910):

you mean 4x-y = 8 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x= -2 y= 6 z= -6

jimthompson5910 (jim_thompson5910):

now double everything in equation 3 to go from x-3y+z= -4 to 2x-6y+2z= -8 Then add this to equation 2 What do you get?

OpenStudy (anonymous):

wait i don't understand double everything, what do you mean

jimthompson5910 (jim_thompson5910):

multiply everything in equation 3 by 2

OpenStudy (anonymous):

ok and i got 2x-6y+2z= -8

jimthompson5910 (jim_thompson5910):

now add this to equation 2

OpenStudy (anonymous):

i got -2x-5y= -4

OpenStudy (anonymous):

what would i do next, cause it looks confusing now

jimthompson5910 (jim_thompson5910):

ok, you now have these 2 equations

jimthompson5910 (jim_thompson5910):

4x-y = 8 -2x-5y = -4

jimthompson5910 (jim_thompson5910):

focus on them only and forget about the rest for now

OpenStudy (anonymous):

ok i got 2x-6y=4

jimthompson5910 (jim_thompson5910):

multiply everything in equation 1 by -5 4x-y = 8 -5*4x+(-5)*(-y) = -5*8 -20x + 5y = -40 Now add -20x + 5y = -40 to -2x-5y = -4

OpenStudy (anonymous):

ok am confused but i got so far is -40x= -44

jimthompson5910 (jim_thompson5910):

-20x + 5y = -40 -2x-5y = -4 ------------------ -22x = -44 -22x = -44 x = -44/(-22) x = 2 Now use this to find y

OpenStudy (anonymous):

ok and then i'd plug it in a equation

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

either -20x + 5y = -40 or -2x-5y = -4 or 4x-y = 8

OpenStudy (anonymous):

ok and am sorry but how did you get 4x-y=8

jimthompson5910 (jim_thompson5910):

it was the result when we added the 1st and 3rd equations (in the very beginning)

OpenStudy (anonymous):

ok thanks i kinda understand it now just some parts idk still kinda like confusing

jimthompson5910 (jim_thompson5910):

go over them again and hopefully they'll make more sense (try to reword them if they still don't)

OpenStudy (anonymous):

ok i will and thanks alot though :) oh y=0, z=-6 right

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (anonymous):

ok thanks